OpenStudy (anonymous):

Wich ordered pair om the form of (a,b) is a solution of this equation? 3a - 4b = 21. The options are [(-3,-2)(-2-3)(7,0)(0,7)]

OpenStudy (anonymous):

I do not understand how it is I am supposed to do this equation

OpenStudy (anonymous):

3(7)-4(0)=21

OpenStudy (anonymous):

You have what it given and then you put the answers given in for 'a' and 'b' to equal 21. so 3(7)-4(0)=? 3*7 is 21 and 4*0 is 0 so 21-0=21 so your answer would be (7,0) but to find the answer you have to plug in the answer choices.

OpenStudy (anonymous):

Okay, that makes since

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Can you medal me? :)

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

No problem.